Solve the system of equations.
x + 5y = 6
2x - 8y = -24
How do I do that?

Found 2 solutions by lynnlo, Vladdroid:
Answer by lynnlo(4176)

(1)x+5y=6
2x-8y-24
-2x-10y=-12
2x-8y=-24
-18y=-36
y=2
x+5(2)=6
x=-4
(2)3x+y=9
2x-4y=6
12x+4y=36
2x-4y=6
14x=42
x=3
3(3)+y=9
y=0

Answer by Vladdroid(91)

(1)x+5y=6
2x-8y-24
-2x-10y=-12
2x-8y=-24
-18y=-36
y=2
x+5(2)=6
x=-4
(2)3x+y=9
2x-4y=6
12x+4y=36
2x-4y=6
14x=42
x=3
3(3)+y=9
y=0